OpenStudy (anonymous):

Find the inverse of f(x)=3x-3 F^-1(y)= ? no decimals

OpenStudy (anonymous):

3x = f(x) + 3 x =( f(x) + 3)/ 3 F^-1(y) = (x +3)/3

OpenStudy (anonymous):

ok?